the dimension of a rectangle are 38 feet by 17 feet. find the the length of a diagonal and the acute angle between it and a short side.?
Found 2 solutions by stanbon, rex12:
Answer by stanbon(75887)

the dimension of a rectangle are 38 feet by 17 feet. find the the length of a diagonal and the acute angle between it and a short side.?
-----------------------------------------------
Draw the rectangle and label the sides.
-----
diagonal = sqrt(38^2+17^2) = 41.63 ft
----
angle = invtan(38/17) = 65.9 degrees
----
Cheers,
Stan H.

Answer by rex12(1)

Draw the rectangle and label the sides.
-----
diagonal = sqrt(38^2+17^2) = 41.63 ft
----
angle = invtan(38/17) = 65.9 degrees
----
Cheers,